# Blood Flow in Elastic Vessel Using PINNH. Zhang, R. Chan and X.C. Tai

### v1

@@ 4.1.1 Environment and Softwares
Our methods are implemented across all experiments using the Python programming language coupled with the Pytorch framework. The Finite Element Method is implemented using the open-source package FEniCS. It’s noteworthy that the computational meshes are generated using Gmsh; however, certain meshes, generated automatically, necessitate manual modifications to make them computable with FEniCS. A Linux server equipped with four Tesla A600 Graphics Processing Units is employed to execute these experiments.
@@ 4.1.2 Hyperparameters
The most important consideration for the weights on the loss functions is ensuring the satisfaction of the boundary and initial conditions. To achieve this, we first train the network to minimize the residuals with $\alpha_{\text{ns}}=0$, $\alpha_{\text{bdr}}^{\text{f}}=1$, and $\alpha_{\text{init}}^{\text{f}}=10^{-1}$ for the fluid problem at the beginning. This setting is motivated by the treatment of initial and boundary conditions in conventional finite element methods, where they are enforced strictly. When the network $N_{u}$ and $N_{p}$ are sufficiently trained, we set $\alpha_{\text{ns}}=10^{-7}$ and update it by $\alpha_{\text{ns}}=10\times\alpha_{\text{ns}}$ until it converges for the current $\alpha_{\text{ns}}$. This dynamic update helps the neural solver to sufficiently satisfy the initial and boundary conditions before optimizing the network to produce accurate solutions for velocity and pressure. Regarding the solid problem, based on empirical experiments, we set $\alpha_{\text{sc}}=1$, $\alpha_{\text{he}}=10$, $\alpha_{\text{bdr}}^{\text{s}}=10^{-2}$, and $\alpha_{\text{init}}^{\text{s}}=10^{-2}$.
We set $M=N=1000$ points with temporal and spatial coordinates sampled on each domain and boundary, as discussed in Section 3.3. Adam optimizer with a learning rate of $1\times 10^{-3}$ is employed to initialize and optimize the learnable variables of neural network. Regarding the network architecture, unless specified otherwise, the network depth is set to $12$ for $N_{u}$, $N_{p}$, and $N_{d}$. The width is set to $20$ for $N_{u}$ and $N_{d}$, and $10$ for $N_{p}$.
For parameters in Algorithm 1, a training unit for the fluid problem contains $m=100$ training epochs. Within each unit, $N_{u}$ is trained for the first $m_{1}=80\%\times m=80$ epochs, while $N_{p}$ is trained for the remaining $m_{2}=20\%\times m=20$ epochs. The training unit for the solid problem also contains $m=100$ training epochs.
For parameters in Algorithm 1, when $\alpha_{\text{ns}}=0$, $N_{u}$ and $N_{p}$ are optimized for a maximum of $m_{\text{fluid}}=20$ units which can be stopped when it converges before that. The convergence parameter is $\epsilon=10^{-2}$. After the initialization stage, we set $\alpha_{\text{ns}}=10^{-7}$, which is then multiplied by $10$ for $E=5$ times. In the phase of iteratively solving the two sub-problems, we run $m_{\text{solid}}=5$ training units to solve the solid problem, followed by a maximum of $m_{\text{fluid}}=20$ training units for the fluid problem. There are a maximum of $F=6$ alternations for solving the FSI problem.
@@ 4.1.3 Error Metric
$E(\boldsymbol{u}^{*},\boldsymbol{u}^{\text{GT}})=t_{l}\sum_{\boldsymbol{t}\in\mathcal{T}}\frac{\sum_{\mathcal{P}\in\mathcal{F}}A(\mathcal{P})\frac{\sum_{\boldsymbol{p}\in\mathcal{P}}|\boldsymbol{u}^{*}(\boldsymbol{p})-\boldsymbol{u}^{\text{GT}}(\boldsymbol{p})|^{2}}{|\mathcal{P}|}}{\sum_{\mathcal{P}\in\mathcal{F}}A(\mathcal{P})\sum_{\boldsymbol{p}\in\mathcal{P}}\frac{|\boldsymbol{u}^{\text{GT}}(\boldsymbol{p})|^{2}}{|\mathcal{P}|}},$ (35)
@@ 4.2 Blood Vessel with Elastic Wall
Figure 4: The illustration for the velocity magnitude (up row) and pressure (bottom row) variation at time points 0.01, 0.21, 0.41, 0.61, and 0.81. The red outside indicates the vessel wall. The color inside the lumen indicates the quantity magnitude. For a more convenient observation, the displacement vectors are enlarged by a factor of 10.
@@
Because of the specific geometry, we can simplify the problem by transforming it from a 3-D coordinate system ($x$, $y$, $z$) to a 2.5-D cylindrical coordinate system ($r$, $\theta$, $z$). This simplification allows us to fix the longitudinal section while varying the radial angle $\theta$, effectively reducing the problem to a 2-D setting. This transformation streamlines the computation and conserves computational resources. It also needs to be mentioned again that the longitudinal direction is designated as the $z$-axis, which is the same as the original coordinate system illustrated in 2.
@@
$\boldsymbol{u}=0\quad\text{ in }\Omega^{\text{f}}_{0},t=0.$ (36)
@@
$\boldsymbol{u}(\boldsymbol{x},t)=g(\boldsymbol{x})(0.1-0.1\cos{(2\pi t)})\quad\text{ on }\Gamma^{\text{ in }}_{t},t\in[0,T].$ (37)
On the outlet of the vessel segment, no stress is imposed. This condition is represented as a neutral boundary condition, written as
$\left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-p\mathbf{I}\right]\cdot\mathbf{n}=\mathbf{0}\quad\text{ on }\ \Gamma_{t}^{\text{out}},t\in[0,T].$ (38)
@@
$\mathbf{u}=\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad\text{ on }\Gamma_{t}^{\text{w}},t\in[0,T].$ (39)
Thus, in the fluid problem, the functional $F^{\text{f}}_{\text{init}}$ that indicates the initial condition and $F^{\text{f}}_{\text{bdr}}$ that indicates the boundary condition can be expressed as
$F^{\text{f}}_{\text{init}}$ $=\boldsymbol{u}\qquad\qquad\qquad\qquad\qquad\qquad\text{ in }\Omega^{\text{f}}_{0},t=0.$ (40)
$F^{\text{f}}_{\text{bdr}}$ =$=\begin{cases}\boldsymbol{u}-g(0.1-0.1\cos{(2\pi t)})\quad&\text{in }\Gamma^{\text{ in }}_{t},t\in[0,T],\\ \left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-p\mathbf{I}\right]\cdot\mathbf{n}\quad&\text{ on }\Gamma_{t}^{\text{out}},t\in[0,T],\\ \mathbf{u}-\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad&\text{ on }\Gamma_{t}^{\text{w}},\,\,t\in[0,T].\end{cases}$ (41)
@@
$\mathbf{d}=\mathbf{0}\quad\text{ in }\Omega_{0},t=0.$ (42)
@@
$\mathbf{d}=\mathbf{0}\quad\text{ on }\partial\Gamma_{0}^{\text{w}},t=0.$ (43)
Therefore, the functionals $F^{\text{s}}_{\text{init}}$ and $F^{\text{s}}_{\text{bdr}}$, indicating the initial and boundary conditions in the solid problem, can be expressed as:
$F^{\text{s}}_{\text{init}}$ $=\boldsymbol{d}\quad\text{ in }\Omega_{0},t=0,$ (44)
$F^{\text{s}}_{\text{bdr}}$ $=\mathbf{d}\quad\text{ on }\partial\Gamma_{0}^{\text{w}},t=0.$ (45)
Figure 5: The temporal profiles of velocity magnitude and pressure at different points. From left to right, the upper three plots illustrate velocity magnitude on the centerline’s inlet, middle, and outlet. The lower three represent pressure magnitude for the same corresponding points. The notations H1, H2, H3, and H4 in the legend denote the grids whose minimum length $h$ are set to be $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$, respectively. T1, T2, and T3 denote time step lengths $1\times 10^{-1}$, $1\times 10^{-2}$, and $1\times 10^{-3}$.
@@
Table 1: The comparison between our methods and FEM. The relative errors for velocity magnitude and pressure are reported. The FEM result using a high-resolution grid with a small time step is taken as the ground truth. In the table, $h^{-1}$ signifies the inverse of the minimum length $h$, used to discretize the geometry. Since our method is mesh-free, we test with different numbers of sample points in the column of the results by our method, which is of different time step lengths for FEM.
With these critical conditions and material properties defined, we are now fully equipped to proceed with solving the simulation problem for blood flow in elastic vessels, as outlined in Section 3. The results of our computations are presented in Table 1 where the relative errors for velocity magnitude and pressure are reported. In this comparison, we utilize the solution computed by the Finite Element Method (FEM) on a grid with a minimum length of $\frac{1}{16}\times 10^{-3}$ meters and a time step of $10^{-3}$ seconds as the ground truth. In the table, $h^{-1}$ signifies the inverse of the minimum length $h$, which is used to discretize the geometry. The results for simulation using FEM also relates to time step size, which is set to be three different lengths in our experiments. Notably, our method is mesh-free, so we don’t need to configure for time step and discretization length. Instead, we test with a different number of sample points, denoted as $N$, $M$, and $K$ as discussed in Section 3.
The results indicate that our method performs impressively with $N$, $M$, and $K$ set to $1000$, achieving high accuracy. Remarkably, the accuracy achieved with these settings is comparable to traditional grid-based methods with a minimum length of $1\times 10^{-4}$ meters and a time step of $10^{-2}$ seconds. However, our approach significantly reduces the computational time.
@@
These observations are further supported by the plots in Figure 5, which depict the temporal profiles of velocity magnitude and pressure at different points. The upper three plots illustrate velocity magnitude, while the lower three represent pressure. To understand the plots, from left to right, they correspond to the points on the centerline’s inlet, middle, and outlet. The symbols H1, H2, H3, and H4 in the legend denote the grids whose minimum length $h$ are set to be $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$, respectively. Similarly, the time step lengths are denoted as T1, T2, and T3. For instance. As an example, the H4T3 corresponds to a grid with a minimum length of $\frac{1}{16}\times 10^{-3}$ meters and a time step of $1\times 10^{-3}$ seconds.
Upon closer inspection, we have noticed that the time steps of $1\times 10^{-1}$ and $1\times 10^{-2}$ can introduce oscillations that lead to more significant errors over time while our approach consistently produces a smooth solution, which is more accurate and usable in practice.
@@ 4.3 Self Ablation
Next, we delve into the influence of different neural network architectures on the model performance. In this section, we maintain the same geometric setup, initial conditions, boundary conditions, and other parameters outlined in the previous simulation problem. The primary variation lies in the neural network architecture, specifically in its depth and the total width for $N_{u}$ and $N_{p}$. In this study, we employ two different width settings. The first setting allocates a width of $20$ to $N_{u}$ and $10$ to $N_{p}$, resulting in a total width of $30$. For the second setting, with a total width of $60$, we maintain the same proportion, assigning $40$ to $N_{u}$ and the remaining $20$ to $N_{p}$. All tests are conducted under identical optimization settings as detailed in Section 4.1.
@@
Another important feature of the proposed methods is its parallelizability. Benefiting from the well-developed PyTorch package, using multiple GPUs to optimize neural networks in a short time is straightforward and simple. Given in Table.3, we reported the time needed when using different numbers of GPUs for training a network with a depth of $12$ and width of $30$.

### v4

@@ 4.1.1 Environment and Softwares
Our methods are implemented across all experiments using the Python programming language coupled with the Pytorch framework. The Finite Element Method is implemented using the open-source package FEniCS and FSI package turtleFSI[7]. It’s noteworthy that the computational meshes are generated using Gmsh; however, certain meshes, generated automatically, necessitate manual modifications to make them computable with FEniCS. A Linux server equipped with four Tesla A600 Graphics Processing Units is employed to execute these experiments.
@@ 4.1.2 Hyperparameters
The most important consideration for the weights on the loss functions is ensuring the satisfaction of the boundary and initial conditions. To achieve this, we first train the network to minimize the residuals with $\alpha_{\text{ns}}=0$, $\alpha_{\text{bdr}}^{\text{f}}=1$, and $\alpha_{\text{init}}^{\text{f}}=10^{-1}$ for the fluid problem at the beginning. This setting is motivated by the treatment of initial and boundary conditions in conventional finite element methods, where they are enforced strictly. When the network $N_{u}$ and $N_{p}$ are sufficiently trained, we set $\alpha_{\text{ns}}=10^{-7}$ and update it by $\alpha_{\text{ns}}=10\times\alpha_{\text{ns}}$ until it converges for the current $\alpha_{\text{ns}}$. This dynamic update helps the neural solver to sufficiently satisfy the initial and boundary conditions before optimizing the network to produce accurate solutions for velocity and pressure. Regarding the solid problem, based on empirical experiments, we set $\alpha_{\text{sc}}=1$, $\alpha_{\text{he}}=10$, $\alpha_{\text{bdr}}^{\text{s}}=10^{-1}$, and $\alpha_{\text{init}}^{\text{s}}=10^{-2}$.
We set $M=N=1000$ points with temporal and spatial coordinates sampled on each domain and boundary, as discussed in Section C. Adam optimizer with a learning rate of $1\times 10^{-3}$ is employed to initialize and optimize the learnable variables of neural network. Regarding the network architecture, unless specified otherwise, the network depth is set to $12$ for $N_{u}$, $N_{p}$, and $N_{d}$. The width is set to $20$ for $N_{u}$ and $N_{d}$, and $10$ for $N_{p}$.
For parameters in Algorithm 1, when $\alpha_{\text{ns}}=0$, $N_{u}$ and $N_{p}$ are optimized for a maximum of $m_{\text{fluid}}=2000$ epochs which can be stopped when it converges before that. The convergence parameter is $\epsilon=10^{-1}$. After the initialization stage, we set $\alpha_{\text{ns}}=10^{-8}$, which is then multiplied by $10$ for $E=5$ times. In the phase of iteratively solving the two sub-problems, we run $m_{\text{solid}}=500$ epochs to solve the solid problem, followed by a maximum of $m_{\text{fluid}}=2000$ training units for the fluid problem. There are a maximum of $F=6$ alternations for solving the FSI problem. For parameters in Algorithm 2, training $m_{1}=80$ epochs for $N_{u}$ will follow a training on $N_{p}$ for $m_{2}=20$ epochs on $N_{p}$.
It’s worth mentioning that the models outlined in Section 2 are adaptable to dimensional forms with units and non-dimensional forms as in [19]. Due to the consideration of applying our methods to real clinical simulations, we retain all units by utilizing the CGS unit system in our implementation. However, it is straightforward to convert the formulation into a non-dimensional form by using the Reynolds number.
@@ 4.1.3 Error Metric
(31) $E(\boldsymbol{u}^{*},\boldsymbol{u}^{\text{GT}})=t_{l}\sum_{\boldsymbol{t}\in\mathcal{T}}\frac{\sum_{\mathcal{P}\in\mathcal{F}}A(\mathcal{P})\frac{\sum_{\boldsymbol{p}\in\mathcal{P}}|\boldsymbol{u}^{*}(\boldsymbol{p})-\boldsymbol{u}^{\text{GT}}(\boldsymbol{p})|^{2}}{|\mathcal{P}|}}{\sum_{\mathcal{P}\in\mathcal{F}}A(\mathcal{P})\sum_{\boldsymbol{p}\in\mathcal{P}}\frac{|\boldsymbol{u}^{\text{GT}}(\boldsymbol{p})|^{2}}{|\mathcal{P}|}},$
@@ 4.2 Blood Vessel with Elastic Wall
Figure 3: Illustration of velocity magnitude (first row), pressure (second row), displacement magnitude (third row) and displacement magnitude for wall (forth row) variations at time points 0.01, 0.21, 0.41, 0.61, and 0.81. The first and the second don’t contain the vessel wall. Deformations of the shape are enlarged by a factor of 10 for better visibility.
@@
(32) $\boldsymbol{u}=0\quad\text{ in }\Omega^{\text{f}}_{0},t=0.$
@@
(33) $\boldsymbol{u}(\boldsymbol{x},t)=g(\boldsymbol{x})(10-10\cos{(2\pi t)})\quad\text{ on }\Gamma^{\text{ in }}_{t},t\in[0,T].$
On the outlet of the vessel segment, no stress is imposed. This condition is represented as a natural boundary condition, written as
(34) $\left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-P\mathbf{I}\right]\cdot\mathbf{n}=\mathbf{0}\quad\text{ on }\ \Gamma_{t}^{\text{out}},t\in[0,T].$
@@
(35) $\mathbf{u}=\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad\text{ on }\Gamma_{t}^{\text{w}},t\in[0,T].$
Thus, in the fluid problem, the function $F^{\text{f}}_{\text{init}}$ that indicates the initial condition and $F^{\text{f}}_{\text{bdr}}$ that indicates the boundary condition can be expressed as
(36) $F^{\text{f}}_{\text{init}}$ $=\boldsymbol{u}\qquad\qquad\qquad\qquad\qquad\qquad\text{ in }\Omega^{\text{f}}_{0},t=0.$
(37) $F^{\text{f}}_{\text{bdr}}$ =$=\begin{cases}\boldsymbol{u}-g(10-10\cos{(2\pi t)})\quad&\text{ in }\Gamma^{\text{ in }}_{t},t\in[0,T],\\ \left[\mu\left(\nabla^{\intercal}\mathbf{u}+\nabla\mathbf{u}\right)-P\mathbf{I}\right]\cdot\mathbf{n}\quad&\text{ on }\Gamma_{t}^{\text{out}},t\in[0,T],\\ \mathbf{u}-\frac{\partial\eta}{\partial t}\mathbf{e}_{r}\quad&\text{ on }\Gamma_{t}^{\text{w}},\,\,t\in[0,T].\end{cases}$
@@
(38) $\mathbf{d}=\mathbf{0}\quad\text{ in }\Omega_{0},t=0.$
@@
(39) $\mathbf{d}=\mathbf{0}\quad\text{ on }\partial\Gamma_{0}^{\text{w}},t=0.$
Therefore, the functions $F^{\text{s}}_{\text{init}}$ and $F^{\text{s}}_{\text{bdr}}$, indicating the initial and boundary conditions in the solid problem, can be expressed as:
(40) $F^{\text{s}}_{\text{init}}$ $=\boldsymbol{d}\quad\text{ in }\Omega_{0},t=0,$
(41) $F^{\text{s}}_{\text{bdr}}$ $=\mathbf{d}\quad\text{ on }\partial\Gamma_{0}^{\text{w}},t=0.$
Figure 4: Temporal profiles of velocity magnitude and pressure at different points. From left to right, the upper three plots illustrate velocity magnitude at the centerline’s inlet, middle, and outlet, while the lower three represent pressure magnitude for the corresponding points. Legend notations: H1, H2, H3, and H4 indicate grids with minimum lengths of $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$ meter, respectively. T1, T2, and T3 denote time step lengths of $1\times 10^{-1}$, $1\times 10^{-2}$, and $1\times 10^{-3}$ second.
@@
Table 1: Comparison between our methods and FEM, reporting relative errors for velocity magnitude and pressure. The FEM result on a high-resolution grid with a small time step is considered ground truth. In the table, $h^{-1}$ signifies the inverse of the minimum length $h$, used to discretize the geometry. Results by our method are shown with different numbers of sample points, corresponding to different time step lengths for FEM.
With these critical conditions and material properties defined, we are now fully equipped to proceed with solving the simulation problem for blood flow in elastic vessels, as outlined in Section 3. The results of our computations are presented in Table 1 where the relative errors for velocity magnitude and pressure are reported. In this comparison, we utilize the solution computed by the Finite Element Method (FEM) on a grid with a minimum length of $\frac{1}{16}\times 10^{-3}$ meter and a time step of $10^{-3}$ seconds as the ground truth. In the table, $h^{-1}$ signifies the inverse of the minimum length $h$, which is used to discretize the geometry. The results for simulation using FEM also relates to time step size, which is set to be three different lengths in our experiments. Notably, our method is mesh-free, so we don’t need to configure for time step and discretization length. Instead, we test with a different number of sample points, denoted as $N$, $M$, and $K$ as discussed in Section 3.
The results indicate that our method performs impressively with $N$, $M$, and $K$ set to $1000$, achieving high accuracy. Remarkably, the accuracy achieved with these settings is comparable to traditional grid-based methods with a minimum length of $1\times 10^{-4}$ meter and a time step of $10^{-2}$ seconds. However, our approach significantly reduces the computational time.
@@
These observations are further supported by the plots in Figure 4, which depict the temporal profiles of velocity magnitude and pressure at different points. The upper three plots illustrate velocity magnitude, while the lower three represent pressure. To understand the plots, from left to right, they correspond to the points on the centerline’s inlet, middle, and outlet. The symbols H1, H2, H3, and H4 in the legend denote the grids whose minimum length $h$ are set to be $\frac{1}{2}\times 10^{-3}$, $\frac{1}{4}\times 10^{-3}$, $\frac{1}{8}\times 10^{-3}$, and $\frac{1}{16}\times 10^{-3}$ meter, respectively. Similarly, the time step lengths are denoted as T1, T2, and T3. For instance. As an example, the H4T3 corresponds to a grid with a minimum length of $\frac{1}{16}\times 10^{-3}$ meter and a time step of $1\times 10^{-3}$ second.
Upon closer inspection, we have noticed that the time steps of $1\times 10^{-1}$ and $1\times 10^{-2}$ seconds can introduce oscillations that lead to more significant errors over time while our approach consistently produces a smooth solution, which is more accurate and usable in practice.
Figure 5: Illustration of loss value variations across training epochs, color-coded to depict changes in the weighting parameter $\alpha_{\text{ns}}$ and alternation between fluid and solid problems. Non-black curves represent fluid problem losses at different $\alpha_{\text{ns}}$ values, with y-axes on the left. Black curves denote solid problem losses, associated y-axes on the right.
The success of the proposed model greatly benefits from the carefully designed optimization strategy, which incorporates dynamic weighting parameters and iterative training for both fluid and solid problems. In Figure 5,we present the variation of the loss values across training for 503 every 100 epochs. The non-black curves represent the loss values for the fluid problem at various $\alpha_{\text{ns}}$ values, whose y-axes are on the left of the plot. The black curves denote the loss values for the solid problem, whose measure is depicted by the right y-axes.
The left y-axes in blue are the loss values for the fluid problem in different $\alpha_{\text{ns}}$ values, while the right y-axes in black are the loss values for the solid problem.
In this plot, distinct segments are color-coded to indicate changes in the weighting parameter $\alpha_{\text{ns}}$. Notably, we observe that as the total loss converges, $\alpha_{\text{ns}}$ is incrementally updated to a higher value. With each increase in $\alpha_{\text{ns}}$, the total loss dramatically changed into a high value, prompting the optimizer to subsequently work on reducing the updated error values. For the matching conditions, the associated residual term experiences a substantial surge in value whenever $\alpha_{\text{ns}}$ is augmented, before being optimized to a lower magnitude. This behavior underscores the efficacy of our dynamic weighting strategy, wherein emphasis is initially placed on improving the matching conditions while concurrently optimizing the Navier-Stokes equation residual.
During the second stage of training, characterized by alternating between the fluid and solid problems (indicated by blue and black segments), the loss steadily diminishes and ultimately converges after four such alternations. Notably, the final training segment for the fluid problem exhibits a marginal reduction in training loss compared to its preceding segment, suggesting a trend toward convergence.
@@ 4.3 Self Ablation
Ours Sigmoid Single NN
No. of Parameters 5473 5473 9513
Velocity Relative Error 0.0128 0.0155 0.0231
Pressure Relative Error 0.0239 0.0286 0.0514
Computational Time 863 1121 1314
Table 3: Comparison of network architectures for fluid problem: Our method uses two FNNs with 12 layers and a width of 30, employing alternating activation functions. “Sigmoid” refers to an architecture with only Sigmoid activations, while “Single NN” indicates a single network with alternating ReLU and Sigmoid activations like ours.
@@
Next, we delve into the influence of different neural network architectures on the model performance. In this section, we maintain the same geometric setup, initial conditions, boundary conditions, and other parameters outlined in the previous simulation problem. The primary variation lies in the neural network architecture.
Network Width and Depth Regarding the depth and the total width for the networks $N_{u}$ and $N_{p}$, we employ two different width settings. The first setting allocates a width of $20$ to $N_{u}$ and $10$ to $N_{p}$, resulting in a total width of $30$. For the second setting, with a total width of $60$, we maintain the same proportion, assigning $40$ to $N_{u}$ and the remaining $20$ to $N_{p}$. All tests are conducted under identical optimization settings as detailed in Section 4.1.
@@
Activation Function: The proposed architecture features an alternating activation function between ReLU and Sigmoid. While the smoothness of Sigmoid is necessary for ensuring at least twice-differentiable solutions, it suffers from vanishing gradient issues away from zero, potentially slowing down convergence. In contrast, ReLU lacks this vanishing gradient problem and is computationally more efficient. By alternating between these two functions, we aim to leverage the benefits of both. To assess this approach, we conducted experiments comparing networks with alternating activation functions to those using only Sigmoid, employing identical configurations in terms of depth (12 layers) and total width (30 neurons). Results, presented in Table 3, demonstrate that the alternating activation function setting yields superior accuracy and significantly reduced computational time.
Single vs. Separate Networks: Existing approaches to solve the Navier-Stokes equation typically employ a single network to predict both pressure and velocity. However, this approach often faced convergence challenges due to the complex internal connections within the hidden layers. Additionally, the Navier-Stokes equation involves distinct operators for velocity and pressure, necessitating tailored approaches for their approximation. As a solution, we employ separate networks for pressure and velocity. This approach provides more flexibility, resulting in improved stability and more precise approximation of the partial differential equations. To validate this, we conducted a comparison experiment on the same simulation problem, contrasting our method with the single-network approach. Results, summarized in Table 3, demonstrate that our method achieves superior accuracy in terms of relative errors for both pressure and velocity. Moreover, the optimization time required for convergence is significantly reduced.
Parallel Training Another important feature of the proposed methods is its parallelizability. Benefiting from the well-developed PyTorch package, using multiple GPUs to optimize neural networks in a short time is straightforward and simple. Given in Table.4, we reported the time needed when using different numbers of GPUs for training a network with a depth of $12$ and width of $30$.